An airplane is flying at 300 m/s( 672mi/h. How much time must elapse before a clock in the airplane and one in the ground differ by 1 sec?
 
 

Grade:12th pass

1 Answers

Arun
5 years ago
Let T denote the elapsed time on the ground, and t the elapsed time on the airplane. Then we have 

T = t/sqrt(1-v^2/c^2). 

We solve for 

T - t = 1s. 

T - T sqrt(1-v^2/c^2) = 1 

T = 1/( 1 - sqrt(1-v^2/c^2) ) 

With v= 10^-5 c, this gives 

T = 2*10^10 seconds 

This is 634 years...